# Decentralized Identifier NSFW Not Safe For Work

### v1

@@ 4 Measuring Starter Pack Use (RQ1)
To briefly explore the topics that these starter packs cover, we inspect each starting group and their members. We employ the Large Language Model (LLM) Mistral (Jiang et al. 2023) (see prompts and details in the Appendix B) for this classification. We prompt the LLM to provide both the focus of the starter pack as well as the list of potential members, often listed in the pack description. For instance a starter pack focused on climate might be explicit about potential members being journalists, politicians, activists, or scientists.
@@ 6.1 Methodology
To better understand the content of the posts, we manually annotate 4,000 (1% of all the posts mentioning starter packs) and identify the 11 most common themes expressed in positive, neutral, and negative posts. We then classify the remaining posts into those 11 themes using the Mistral LLM (see Appendix B.2 for the methodology). Note, we include an “other” category for all remaining posts that do not fall into one of the 11 themes.
@@ 8 Conclusion & Future Work
Appendix A Ethical Checklist
A.1 Paper Checklist to be included in your paper
  1. 1.
    For most authors…
    1. (a)
      Would answering this research question advance science without violating social contracts, such as violating privacy norms, perpetuating unfair profiling, exacerbating the socio-economic divide, or implying disrespect to societies or cultures? Yes
    2. (b)
      Do your main claims in the abstract and introduction accurately reflect the paper’s contributions and scope? Yes
    3. (c)
      Do you clarify how the proposed methodological approach is appropriate for the claims made? Yes
    4. (d)
      Do you clarify what are possible artifacts in the data used, given population-specific distributions? Yes
    5. (e)
      Did you describe the limitations of your work? Yes
    6. (f)
      Did you discuss any potential negative societal impacts of your work? NA
    7. (g)
      Did you discuss any potential misuse of your work? NA
    8. (h)
      Did you describe steps taken to prevent or mitigate potential negative outcomes of the research, such as data and model documentation, data anonymization, responsible release, access control, and the reproducibility of findings? NA. Our research relies on public data only.
    9. (i)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? Yes
  2. 2.
    Additionally, if your study involves hypotheses testing…
    1. (a)
      Did you clearly state the assumptions underlying all theoretical results? Yes
    2. (b)
      Have you provided justifications for all theoretical results? Yes
    3. (c)
      Did you discuss competing hypotheses or theories that might challenge or complement your theoretical results? Yes
    4. (d)
      Have you considered alternative mechanisms or explanations that might account for the same outcomes observed in your study? Yes
    5. (e)
      Did you address potential biases or limitations in your theoretical framework? Yes
    6. (f)
      Have you related your theoretical results to the existing literature in social science? Yes
    7. (g)
      Did you discuss the implications of your theoretical results for policy, practice, or further research in the social science domain? Yes
  3. 3.
    Additionally, if you ran machine learning experiments…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? Yes. We provide all the prompts in the appendix. We will release the code after the review process to avoid violating double-blind rules.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? Yes, see the appendix
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? NA
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? Yes
    5. (e)
      Do you justify how the proposed evaluation is sufficient and appropriate to the claims made? Yes
    6. (f)
      Do you discuss what is “the cost“ of misclassification and fault (in)tolerance? NA
Appendix B Appendix: Methodology for Post Classification
@@
B.1 Starter Packs Classification Prompt
@@ 81. Provide two classifications:
B.2 Posts Classification Prompt
@@
B.3 Prompt Template for Queries

### v2

@@ 4 Measuring Starter Pack Use (RQ1)
To briefly explore the topics that these starter packs cover, we randomly sample and inspect 25% of the 58,126 starter packs that include descriptions. We employ the Large Language Model (LLM) Mistral (Jiang et al. 2023) (see prompts and details in the Appendix A) for this classification. We prompt the LLM to provide both the focus of the starter pack as well as the list of potential members, often listed in the pack description. For instance a starter pack focused on climate might be explicit about potential members being journalists, politicians, activists, or scientists.
@@ 6.1 Methodology
To better understand the content of the posts, we manually annotate 4,000 (1% of all the posts mentioning starter packs) and identify the 11 most common themes expressed in positive, neutral, and negative posts. We then classify the remaining posts into those 11 themes using the Mistral LLM (see Appendix A.2 for the methodology). Note, we include an “other” category for all remaining posts that do not fall into one of the 11 themes.
@@ 8 Conclusion & Future Work
Appendix A Appendix: Methodology for Post Classification
@@
A.1 Starter Packs Classification Prompt
@@ 81. Provide two classifications:
A.2 Posts Classification Prompt
@@
A.3 Prompt Template for Queries